# Reconstruction of Dark Matter and Baryon Density From Galaxies: A Comparison of Linear, Halo Model and Machine Learning-Based Methods

### v1

@@ Abstract
For many analyses in cosmology it is necessary to reconstruct the likely distribution of unobserved fields, such as dark matter or baryons, from observed luminous tracers. The dominant approach in cosmology has been to use the so-called halo model, which assumes radially symmetric profiles centered around luminous tracers such as galaxies. More recently, field-level machine learning methods have been proposed that can learn to estimate the unobserved field after being trained on simulations. However, it is unclear whether machine learning methods indeed significantly improve over linear methods or the halo model. In this paper we make a systematic comparison of different approaches to reconstruct dark matter and baryons from galaxy data using the CAMELS simulations. We find the best results using a combined GNN-CNN approach. We also provide a general analysis and visualization of the relationship of matter, baryons, halos and galaxies in these simulations to interpret our results.
@@ 1 Introduction
Within the CDM paradigm, dark matter provides the scaffolding on which structure forms, and baryonic matter, which we will refer to as ‘gas’, provides the medium from which astrophysical objects like stars and galaxies form and evolve. Dark matter, and the majority of gas, is not directly observable. Instead, we observe luminous tracers, e.g. starlight from galaxies or thermal radiation from hot gas, that are embedded within the underlying non-luminous dark matter and gas. It is possible to observe the interaction of photons with dark matter and gas, e.g. through weak gravitational lensing or the kinetic Sunyaev Zel’dovich effect, but this will not be the focus here. Instead, we are concerned with how luminous tracers can be used to make field-level inferences about the distribution of dark matter and gas. Working at the field-level, as opposed to focusing on summary statistics like power spectra, has the advantage of preserving the maximum amount of cosmological information (see e.g. Ref. [1]). Some applications of field-level inference include: incorporating non-Gaussian information in cosmological parameter constraints, incorporating the effect of survey geometry and systematics, testing the environmental-dependence of galaxy formation, and creating templates for cross-correlation studies.
Making field-level inferences requires a model for the connection between luminous tracers and the underlying dark matter and gas. A dominant paradigm for understanding this connection is the halo model of large-scale structure (see e.g. Refs. [2, 3]). Within this construction, all dark matter is assumed to be bound into halos whose positions are correlated by the long-wavelength density field. Each halo is populated with gas or galaxies according to a set of theoretical prescriptions, with parameters calibrated by matching a variety of observations. For galaxies this is referred to as the ‘Galaxy-Halo Connection’ (see e.g. Ref. [4] for a review) and the (probabilistic) set of rules for populating halos with galaxies is the Halo Occupation Distribution (HOD). The HOD depends on the properties of the galaxy sample and dark matter mass of the halo, but might also depend on more complex physics, e.g. the assembly history (a recent assessment can be found here [5]). For gas, a simple model could be a parameterized spherically-symmetric density profile that depends on halo mass (a recent example can be found in Ref. [6]). This profile may be derived based on physical principles, e.g. hydrostatic equilibrium, or measured from hydrodynamic simulations or observations. For computing summary statistics, the halo model provides a physically-motivated description that can be fit to data in order to make inferences about cosmology, galaxies, and gas. However, there are limitations to this approach such as the non-conservation of mass and a poor description of intermediate scales (the halo model can be modified to mitigate these shortcomings, e.g. [7, 8]).
It is also possible to apply the halo model at the field-level. The simplest example of this would be to create a linear transfer function in Fourier space that relates a tracer to the underlying dark matter distribution (see e.g. Ref. [9] for a recent example). The filter leading to the minimum variance estimate of the dark matter field is the cross-power spectrum between the tracer and dark matter, normalized by the auto-spectrum of the tracer, both computed using the halo model (for an example see e.g. Ref. [10]). On linear scales, this corresponds to dividing the tracer field by its linear bias with respect to dark matter. On smaller scales, non-linear effects – such as baryonic feedback and non-linear gravitational evolution – become significant, causing the linear transfer function approach to break down. Another technique that can be applied at the field level is halo abundance matching (HAM) [11, 12, 13]. This technique relates a halo quantity (eg. halo mass ) to a galaxy quantity (eg. stellar mass ) by assuming a monotonic relation and matching the cumulative abundances (). One can then assign dark matter masses to the position of each galaxy to create an estimate of the density field. A downside is that HAM is limited to using only a single galaxy property to infer the host halo characteristics. The HAM approach can be generalized to include more information about the distribution of density in dark matter halos. For example, instead of assigning a mass to the position of the halo, one could ‘paint’ on the associated dark matter density profile, e.g. the Navarro-Frank-White (NFW) profile.
The linear transfer function and HAM approaches described above rely on linear correlations (between fields and galaxy - halo properties, respectively). Non-linear methods can in principle capture more cosmological information, improving our inference of the distribution of dark matter from luminous tracers. One approach that has been extensively explored in the literature is to forward-model galaxy surveys, sampling Gaussian initial conditions, evolving them with N-body simulations, creating mock galaxy surveys and comparing the mocks with data to find the most likely underlying dark matter distribution [14, 15, 16, 17]. This approach is based on causal inferences (because samples are generated by evolving initial conditions with physical equations of motion) rather than linear correlation, and the statistical distribution over initial conditions is Gaussian, allowing precisely defined error bars. However, sampling a large range of scales is computationally challenging. We do not pursue this class of methods further in this work. Another set of approaches relies on machine learning (ML) methods. ML offers a powerful framework for handling high-dimensional problems and capturing complex relationships across a range of scales. For example, Refs. [18, 19, 20, 21, 22, 23] use ML to improve on HAM by predicting halo masses from a variety of galaxy properties in addition to stellar mass. At the field level, Refs. [24, 25] employ diffusion models to reconstruct 2D dark matter fields from galaxy distributions. These approaches can be far more computationally efficient than full forward-modeling, but rely on correlation (no explicit causal inference), and error bars are not rigorously defined from Gaussian initial conditions.
@@
At a basic level, the problem we wish to solve is to find the statistically ideal non-linear map between a point cloud (galaxies) and a pixelated continuous field (dark matter or gas). A class of ML architectures well-suited to point clouds are graph neural networks (GNN); see e.g. [26]. The GNN architecture encodes galaxy features in nodes, and relational information (e.g. relative distance) in vertices connecting them. GNNs have successfully been used to infer the mass of individual halos from galaxy features e.g. [27, 28] or cosmological parameters e.g. [29]. A class of ML architectures well-suited to pixelated continuous fields, a type of regular grid-structured data, are convolutional neural networks (CNN). CNNs have been applied to a wide variety of problems in cosmology e.g. Refs. [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. In Ref. [41], we introduced a hybrid achitecture in which a GNN step is used to map galaxies to dark matter halos and a CNN step is used to map the halo distribution to an inference of the dark matter field in three dimensions. Training the network on the CAMELS suite of cosmological simulations [42, 43, 44, 45], we demonstrated that such a GNN-CNN hybrid architecture could outperform the linear transfer function or a gridded halo mass assignment.
@@
In this paper, we undertake a detailed study of the capabilities of a GNN-CNN hybrid architecture to reconstruct the dark matter field from galaxies, comparing the fidelity of the reconstruction to a variety of other techniques based on the halo model. We provide a detailed visualization of the reconstructed dark matter field using different techniques, highlighting the features that are captured and those that are not. We demonstrate that the GNN-CNN reconstruction trained on CAMELS is superior to halo-model based approaches, and investigate its robustness with respect to unknown feedback parameters. Finally, we comment on cosmological and astrophysical parameter inference using this reconstruction.
@@ 2 Overview of Approaches to Reconstruct Unobserved Fields
The goal of this paper is to reconstruct the unobserved matter distribution and the unobserved baryon distribution from the observed galaxy density . In this section, we first give a high-level overview of several possible reconstruction methods.
Linear transfer function.
@@
as studied for example in [46] in the context of mapping N-body to hydrodynamic simulations, which would make the power spectrum correct by definition.
Of course, a linear transfer function can never improve the cross-correlation coefficient between the observed field and the target field. We also note that a linear transfer function can always be added to other methods, to enforce that the reconstructed field has any desired power spectrum. Note however, that one does not always want the power spectrum of the reconstructed field to be equal to the target field. Often, the output should be inverse covariance weighted so that noisy (small-scale) modes are suppressed. In the above transfer function Eq. (2.2), this will happen if includes the galaxy shot noise.
@@
Point cloud-to-field level machine learning.
In this paper, we further explore the performance of the combined GNN-CNN method from Ref. [41] to directly learn the non-linear galaxy point cloud-to-mass/baryon density field map. We train this architecture on the CAMELS suite of simulations, and compare the fidelity of the reconstruction to that obtained from the linear transfer function and halo painting methods described above. We demonstrate that this ML-based method yields the best reconstruction of the density fields among the methods we explore.
Relation to forward modeling.
Many other works are concerned with forward modeling frameworks to connect matter to luminous matter. In forward modeling, one typically evolves initial conditions to a final matter field by a deterministic forward model (such as a gravitational N-body simulation) which depends on unknown forward model parameters (such as the matter density parameter ):
@@ 4.1 Performance Metric
where is the power spectrum of the field and is the power spectrum of the estimate of that field. The cross-correlation coefficient is the relevant performance metric if the goal is to increase the signal-to-noise ratio of a cross-correlation-based data analysis. We compute the power spectra using Pylians [56]. We do not deconvolve the window functions when plotting the power spectra, since they cannot be removed for the CMD fields. This means that there is a suppression of power at small-scales in the spectra plots.
@@ 7 Robustness Against Unknown Feedback Parameters
We now discuss the robustness of the machine-learned dark matter field to changes in the astrophysical feedback parameters. In all previous discussions, the model was trained and tested on simulations from the same set with the same fiducial values of cosmological and astrophysical parameters. However, it is interesting to see how the model performs on a simulation that has significantly different fiducial parameter values. For this purpose, we use the EX subset of CAMELS-IllustrisTNG. This subset has 4 simulations, one with fiducial values and 3 with the values of the astrophysical parameters at their extremes. We consider three cases:
@@
For this test, we again only evaluate the performance of GNN-CNN and linear transfer function methods. The GNN-CNN is trained on a subset of the CV simulations. For the linear transfer function, the ‘training’ step involves applying a filter (Eq. (2.2)) derived from one of the CV simulations to the EX simulations. In Fig. (17), the cross-correlation coefficients for the GNN-CNN and linear filter fields do not change significantly between the simulations. If there are errors with the individual halo profiles, these get diluted by averaging across all the halos in the simulation. So, even for the most extreme case where there is no feedback, the cross-correlation is not significantly impacted because there are so many halos.
To quantify how biased the results are, we plot where is the GNN-CNN or linear transfer function result. This is shown in Fig. (18). For both the GNN-CNN and linear transfer function methods, the bias generally remains scale-independent. The main exception is the GNN-CNN no feedback reconstruction, which also exhibits the largest bias. This significant drop in performance for the no feedback case is likely due to the extreme change in galaxy number density; this simulation contains over three times the number of galaxies present in the training set. It is promising that the very efficient AGN simulation does not result in a significant bias. Additionally, we anticipate that the GNN-CNN’s performance could improve when applied to different simulations if trained (as in the initial paper [41]) on simulations with varying cosmological and astrophysical parameters.
@@ 9 Conclusions and Outlook
In this paper, we compared different methods of reconstructing continuous fields, in particular dark matter, from discrete tracers. We compared linear transfer function methods (which includes painting uniform profiles on each object), individual halo painting (depending on the halo mass), and finally field-level machine learning. As a baseline, we used a simple stellar mass-weighted grid assignment of galaxies. Our findings are:
@@
In summary, the GNN-CNN method showed a clear improvement over the linear filter technique, which is currently employed in applications such as large-scale velocity reconstruction with kSZ [63, 10, 64, 65, 66, 67, 68, 69, 70] and searches for light bosons using resonant conversion [78, 79, 80]. This indicates that adopting the GNN-CNN setup could increase the quality of these and similar analyses. More broadly, this technique’s ability to provide a field-level reconstruction of dark matter — the fundamental scaffolding of the universe’s matter content — makes it generally applicable to cross-correlation studies, with probes like cosmic shear, 21-cm intensity maps, fast radio bursts, and gravitational wave sources. We explore these and other applications in future work.

### v2

@@ Abstract
For many analyses in cosmology it is necessary to reconstruct the likely distribution of unobserved fields, such as dark matter or non-luminous baryons, from observed luminous tracers. The dominant approach in cosmology has been to use the so-called halo model, which assumes radially symmetric profiles centered around luminous tracers such as galaxies. More recently, field-level machine learning methods have been proposed that can learn to estimate the unobserved field after being trained on simulations. However, it is unclear whether machine learning methods indeed significantly improve over linear methods or the halo model. In this paper we make a systematic comparison of different approaches to reconstruct dark matter and non-luminous baryons, from galaxy data using the CAMELS simulations. These simulations are in a box, allowing us to compare performance on the mildly non-linear scales down to the size of individual halos. We find the best results using a combined GNN-CNN approach. We also provide a general analysis and visualization of the relationship of matter, non-luminous baryons, halos, and galaxies in these simulations to interpret our results.
[…]
@@ 1 Introduction
Within the CDM paradigm, dark matter provides the scaffolding on which structure forms, while baryonic matter, which we will refer to as ‘gas’, provides the medium from which astrophysical objects like stars and galaxies form. Dark matter, and the majority of gas, is not directly observable. Instead, we observe luminous tracers, e.g. starlight from galaxies or thermal radiation from hot gas, that are embedded within the underlying non-luminous dark matter and gas. In this paper we study how luminous tracers can be used to make field-level inferences about the distribution of dark matter and gas, using both classical and machine learning methods. Making field-level inferences requires a model for the connection between luminous observable tracers and the underlying unobservable dark matter and gas fields provided by simulations. These models can be analytical, empirical, or numerical. In all cases, the observable input is a galaxy catalog (real or simulated), and the output is an estimate of the continuous dark matter or gas field.
@@
At a basic level, the problem we wish to solve is to find the statistically ideal non-linear map between a point cloud (galaxies) and a pixelated continuous field (dark matter or gas). While we will also evaluate traditional halo and transfer function based approaches, our focus will be on machine learning. A class of machine learning (ML) architectures well-suited to point clouds are graph neural networks (GNN); see e.g. [1]. The GNN architecture encodes galaxy features in nodes, and relational information (e.g. relative distance) in vertices connecting them. GNNs have successfully been used to infer the mass of individual halos from galaxy features e.g. [2, 3] or cosmological parameters e.g. [4]. A class of ML architectures well-suited to pixelated continuous fields, a type of regular grid-structured data, are convolutional neural networks (CNN). CNNs have been applied to a wide variety of problems in cosmology e.g. Refs. [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. In Ref. [16], we introduced a hybrid GNN-CNN architecture that maps a galaxy catalog directly to the dark matter field in three dimensions. Training the network on the CAMELS suite of cosmological simulations [17, 18, 19, 20], we demonstrated that such a GNN-CNN hybrid architecture could outperform the linear transfer function or a gridded halo mass assignment.
@@
In this paper, we undertake a detailed study of the capabilities of a GNN-CNN hybrid architecture to reconstruct the dark matter field from galaxies on the scales of 25 Mpc/h, down to the scale of individual halo, comparing the fidelity of the reconstruction to a variety of other techniques based on the halo model. In Section (2), we provide an overview of techniques from the literature to reconstruct unobserved fields. Throughout this paper, we use simulations for the galaxies and dark matter fields so we can benchmark against a ground truth dark matter, these simulations are described in Section (3). The ML and traditional methods we use to reconstruct the unobserved fields are outlined in Section (4). In the results, Section (5), we provide a detailed visualization of the reconstructed dark matter field using different techniques, highlighting the features that are captured and those that are not. We demonstrate that the GNN-CNN reconstruction trained on CAMELS is superior to halo-model based approaches. This is also shown for the reconstructed gas field in Section (6) and we investigate the GNN-CNN’s robustness with respect to unknown feedback parameters in Section (7). We comment on cosmological and astrophysical parameter inference using this reconstruction in Section (8). In Appendix (A), we also discuss explicitly how baryonic uncertainty can be taken into account by marginalizing over a bias parameter, in the specific case of kinetic Sunyaev-Zeldovich velocity reconstruction, which is a major physical application for our method. Finally, we summarize our results in Section (9) and discuss future avenues to apply this to real data.
@@ 2 Overview of Approaches to Reconstruct Unobserved Fields
The goal of this paper is to reconstruct the unobserved matter distribution and the unobserved baryon distribution from the observed galaxy density . We assess the performance of different reconstruction methods using simulations, with the future goal to apply the best method to galaxy observations. In this section, we first give a high-level overview of several possible reconstruction methods.
2.1 Linear Transfer Function
@@
as studied for example in [21] in the context of mapping N-body to hydrodynamic simulations, which would make the power spectrum correct by definition. A further recent example of the use of a transfer function to model the gas distribution can be found in Ref. [22]. The transfer function can either be fit directly to simulations or calculated using the halo model - which is introduced in the following Section (2.2).
Of course, a linear transfer function does not improve the cross-correlation coefficient between the observed field and the target field, which is invariant under this re-scaling. We also note that a linear transfer function can always be added to other methods, to enforce that the reconstructed field has any desired power spectrum. Note however, that one does not always want the power spectrum of the reconstructed field to be equal to the target field. Often, the output should be inverse covariance weighted so that noisy (small-scale) modes are suppressed. In the above transfer function Eq. (2.2), this will happen if includes the galaxy shot noise.
2.2 Dark Matter and Gas Reconstruction Using Halo Models
A dominant paradigm for understanding the connection between luminous tracers and the dark matter and gas distribution is the halo model of large-scale structure (see e.g. Refs. [23, 24]). Within this analytic construction, all dark matter is assumed to be bound into halos whose positions are correlated by the long-wavelength density field. Each halo is populated with gas or galaxies according to a set of theoretical prescriptions, with parameters calibrated by matching a variety of observations. For galaxies this is referred to as the ‘Galaxy-Halo Connection’ (see e.g. Ref. [25] for a review) and the (probabilistic) set of rules for populating halos with galaxies is the Halo Occupation Distribution (HOD). The HOD depends on the properties of the galaxy sample and dark matter mass of the halo, but might also depend on more complex physics, e.g. the assembly history (a recent assessment can be found here [26]). For gas, a simple model could be a parameterized spherically-symmetric density profile that depends on halo mass (a recent example can be found in Ref. [27]). This profile may be derived based on physical principles, e.g. hydrostatic equilibrium, or measured from hydrodynamic simulations or observations. For computing summary statistics, the halo model provides a physically-motivated description that can be fit to data in order to make inferences about cosmology, galaxies, and gas. However, there are limitations to this approach such as the non-conservation of mass and a poor description of intermediate scales (the halo model can be modified to mitigate these shortcomings, e.g. [28, 29]).
@@ 2.2 Dark Matter and Gas Reconstruction Using Halo Models
2.3 Dark Matter and Gas Reconstruction Using Machine Learning
Another set of approaches relies on ML methods. These are trained on simulations and can then be applied to galaxy observations. ML offers a powerful framework for handling high-dimensional problems and capturing complex relationships across a range of scales. For example, Refs. [34, 35, 36, 37, 38, 39] use ML to improve on halo abundance matching by predicting halo masses from a variety of galaxy properties in addition to stellar mass. At the field level, Refs. [40, 41] employ diffusion models to reconstruct 2D dark matter fields from galaxy distributions, and Refs. [42, 43] use CNNs to reconstruct dark matter density, velocity, and tidal fields from halo/galaxy samples.
Galaxy catalogs are better represented on a graph than a grid - to capture their discreteness and the tabular nature of their observed properties. This structure is well captured by GNNs. In contrast, the target dark matter and gas fields are continuous fields defined on regular grids, which makes CNNs an appropriate choice. Therefore, a suitable framework for field level reconstruction is to combine the above ML methods and have a hybrid GNN-CNN model [16]. In this paper, we further explore the performance of the combined GNN-CNN method from Ref. [16] to directly learn the non-linear galaxy point cloud-to-mass/baryon density field map. We train this architecture on the CAMELS suite of simulations, and compare the fidelity of the reconstruction to that obtained from the linear transfer function and halo painting methods described above. We demonstrate that this ML-based method yields the best reconstruction of the density fields among the methods we explore.
2.4 Relation to Forward Modeling
One numerical approach that has been extensively explored in the literature is to forward-model galaxy surveys, sampling Gaussian initial conditions, evolving them with N-body simulations, creating mock galaxy surveys and comparing the mocks with galaxy data to find the most likely underlying dark matter distribution [44, 45, 46, 47]. This approach is based on causal inferences (because samples are generated by evolving initial conditions with physical equations of motion) rather than linear correlation, and the statistical distribution over initial conditions is Gaussian, allowing precisely defined error bars.
Typically, forward modeling works by evolving initial conditions to a final matter field by a deterministic forward model (such as a gravitational N-body simulation) which depends on unknown forward model parameters (such as the matter density parameter ):
@@ 4.1 Performance Metric
where is the power spectrum of the field, is the power spectrum of the estimate of that field, and is a cross-spectrum between the two. The cross-correlation coefficient is the relevant performance metric if the goal is to increase the signal-to-noise ratio of a cross-correlation-based data analysis. We compute the power spectra using Pylians [53]. We do not deconvolve the window functions when plotting the power spectra, since they cannot be removed for the CMD fields. This means that there is a suppression of power at small-scales in the spectra plots.
@@ 7 Robustness Against Unknown Feedback Parameters
We now discuss the robustness of the machine-learned dark matter field to changes in the astrophysical feedback parameters. In all previous discussions, the model was trained and tested on simulations from the same set with the same fiducial values of cosmological and astrophysical parameters. We now relax this assumption and discuss generalization outside of the training data. In our previous work, Ref. [16], we explored the sensitivity of the GNN-CNN to training on simulations with varied cosmological and astrophysical parameters. We found that the cross-correlation coefficients were unchanged whether it was trained conditionally or unconditionally on the parameter values. In this paper however, we focus only on varying astrophysical parameters because we have precise measurements of cosmological parameters from data, whereas astrophysics is more uncertain - recent kSZ data shows that feedback is much stronger than our simulations predict e.g. [74]. Therefore, in the following we test how the reconstruction results are biased if the fiducial astrophysics differs from reality.
Performance on simulations with different astrophysical parameters.
It is interesting to see how the model performs on a simulation that has significantly different fiducial astrophysical parameter values. For this purpose, we use the EX subset of CAMELS-IllustrisTNG. This subset has 4 simulations, one with fiducial values and 3 with the values of the astrophysical parameters at their extremes (but fixed cosmological parameters). We consider three cases:
@@
For this test, we again only evaluate the performance of GNN-CNN and linear transfer function methods. The GNN-CNN is trained on a subset of the CV simulations. For the linear transfer function, the ‘training’ step involves applying a filter (Eq. (2.2)) derived from one of the CV simulations to the EX simulations. In Fig. (17), the cross-correlation coefficients for the GNN-CNN and linear filter fields change only minimally between the simulations. If there are errors with the individual halo profiles, these get diluted by averaging across all the halos in the simulation. So, even for the most extreme case where there is no feedback, the cross-correlation is not strongly impacted because there are so many halos.
To quantify how biased the results are, we plot where is the GNN-CNN or linear transfer function result. This is shown in Fig. (18). For both the GNN-CNN and linear transfer function methods, the bias generally remains scale-independent. The main exception is the GNN-CNN no feedback reconstruction, which also exhibits the largest bias. This significant drop in performance for the no feedback case is likely due to the extreme change in galaxy number density; this simulation contains over three times the number of galaxies present in the training set. It is promising that the very efficient AGN simulation does not result in a significant bias. Additionally, we anticipate that the GNN-CNN’s performance could improve when applied to different simulations if trained on simulations with varying cosmological and astrophysical parameters. We briefly explored in our prior work, [16], but only assessed the cross-correlation coefficients not the bias.
Robust cosmological constraints with squeezed limit observables. One may ask whether a ML model trained on simulations, which will have incorrect baryonic physics compared to reality, can ever be used in practice for cosmological parameter analysis. A clean case where such baryonic uncertainty can be marginalized over, and one can thus achieve robust results, are squeezed limit observables. A recent example for using ML in a squeezed limit analysis is [75, 76], where a neural network was used to estimate in a way that is robust to baryonic uncertainty. This is possible because small-scale baryonic physics results in a constant multiplicative bias parameter on large scales that can be marginalized over when estimating cosmological parameters. Here, we are proposing a different application of ML in the squeezed limit. We can use our GNN-CNN method to reconstruct the electron density from observed galaxies, and then use this estimate as an input field to the quadratic kSZ velocity reconstruction estimator. As we recall in Appendix (A), a mismatch between the true (in reality) and assumed (in simulations or theoretical modeling) electron distribution results in a multiplicative bias of the reconstructed velocity. This bias can easily be marginalized over when obtaining cosmological parameter constraints from kSZ velocity reconstruction. Nevertheless, as we showed in our previous work [70] and recall in Appendix (A), a better electron reconstruction raises the overall signal-to-noise in the analysis, and thus tightens cosmological constraints. We can thus robustly apply our GNN-CNN method in this setup, and obtain unbiased cosmological constraints despite baryonic uncertainty in the training data of our GNN-CNN. This will be explored in more detail in future work. We note however that this robustness only holds for squeezed limit observables. In the next section we briefly comment on more general analyses.
@@ 9 Conclusions and Outlook
In this paper, we compared different methods of reconstructing continuous fields, in particular dark matter, from discrete tracers. We compared linear transfer function methods (which includes painting uniform profiles on each object), individual halo painting (depending on the halo mass), and finally field-level ML. As a baseline, we used a simple stellar mass-weighted grid assignment of galaxies. Our findings are:
@@
In summary, the GNN-CNN method showed a clear improvement over the linear filter technique, which is currently employed in applications such as large-scale velocity reconstruction with kSZ [60, 61, 62, 63, 64, 65, 66, 67, 68] and searches for light bosons using resonant conversion [79, 80, 81]. This indicates that adopting the GNN-CNN setup could increase the quality of these and similar analyses. We show this explicitly for kSZ velocity reconstruction in Appendix (A), and also discuss how to marginalize over baryonic uncertainty in this case. More broadly, this technique’s ability to provide a field-level reconstruction of dark matter — the fundamental scaffolding of the universe’s matter content — makes it generally applicable to cross-correlation studies, with probes like cosmic shear, 21-cm intensity maps, fast radio bursts, and gravitational wave sources. We explore these and other applications in future work.